# Towards Unifying Evaluation of Counterfactual Explanations: Leveraging Large Language Models for Human-Centric Assessments

### v2

@@ Abstract
As machine learning models evolve, maintaining transparency demands more human-centric explainable AI techniques. Counterfactual explanations, with roots in human reasoning, identify the minimal input changes needed to obtain a given output and, hence, are crucial for supporting decision-making. Despite their importance, the evaluation of these explanations often lacks grounding in user studies and remains fragmented, with existing metrics not fully capturing human perspectives. To address this challenge, we developed a diverse set of 30 counterfactual scenarios and collected ratings across 8 evaluation metrics from 206 respondents. Subsequently, we fine-tuned different Large Language Models (LLMs) to predict average or individual human judgment across these metrics. Our methodology allowed LLMs to achieve an accuracy of up to 63% in zero-shot evaluations and 85% (over a 3-classes prediction) with fine-tuning across all metrics. The fine-tuned models predicting human ratings offer better comparability and scalability in evaluating different counterfactual explanation frameworks.
@@ Introduction
The rapid adoption of AI across various domains has significantly increased the urgency for explainable AI models. Counterfactual explanations, which address the question ”How should the input be different in order to change the model’s decision outcome?” (Wachter, Mittelstadt, and Russell 2017), not only clarify the machine’s reasoning but also suggest potential changes that users might implement to achieve different results. These explanations enhance user trust and understanding by providing a richer mental representation compared to causal explanations (Warren, Byrne, and Keane 2023). Additionally, counterfactual explanations align closely with human cognitive processes (Miller 2019), as they provide alternative hypothetical realities that are pervasive in our natural reasoning (Byrne 2002).
@@
In light of these considerations, this paper addresses the following question: Can the evaluation process of counterfactual explanations be effectively automated using LLMs? To answer this question, we created a diverse set of 30 counterfactual scenarios that were designed to vary across multiple dimensions of explanatory qualities. The scenarios were evaluated by 206 human respondents in overall satisfaction and metrics of feasibility, consistency, completeness, trust, fairness, complexity and understandability. Next, we divided data for fine-tuning several LLM models to assess every metric score and compared the results to human data on a reserved test set. The pipeline can be seen in Figure 1.
@@
  • First, we present a diverse dataset of human-evaluated counterfactual explanations, encompassing a variety of metrics and scenarios, which could serve both for benchmarking and for training better causal representations of data, as demonstrated in (Chen et al. 2023).
@@
In recent years, some user studies have been conducted with tabular counterfactual data. For instance, (Warren, Byrne, and Keane 2023) conducted a study with 127 participants to compare the effects of counterfactual and causal explanations on objective prediction accuracy and subjective judgments of satisfaction and trust. (Bove et al. 2023) explored the impact of plural counterfactual examples on objective understanding and a modified Explanation Satisfaction Scale (Hoffman et al. 2018) in a lab study with 112 participants. (Förster et al. 2021) conducted a study with 46 participants assessing explanation realism and typicality. Two user studies have benchmarked counterfactual methods for perceived practicality of users in a study with 135 participants (Ghazimatin et al. 2020), and an online study with 500 responders (Spreitzer, Haned, and van der Linden 2022). Additionally, (Akula et al. 2022) tested their approach on image data, evaluating justified trust as quantitative metric and explanation satisfaction as qualitative metric.
@@
Predicting human evaluation with machine learning has gained widespread acceptance in various domains, such as human-computer interaction (Kiseleva et al. 2016), recommendation systems (Siro, Aliannejadi, and De Rijke 2023), speech quality assessment (Reddy, Gopal, and Cutler 2022), etc. The progressive advancement of LLMs’ causal reasoning abilities (Bhattacharjee et al. 2024) suggests their usage in the context of explainability, as their natural language explanations exhibit qualities similar to human output (Castelnovo et al. 2024) and the explanatory process can be further enhanced through a post-output chat pipeline (Slack et al. 2023). LLMs have also been used to evaluate and model user satisfaction, providing insights into choices and preferences (Kim et al. 2024), and as artificial user-model tuning pairs (Gao et al. 2024).
@@
To select the dimensions for our study, we reviewed literature on qualitative metrics that influence human judgments. Among the most frequently cited explanatory virtues are coherence and simplicity (Mackonis 2013), aligning with the understanding of human mental models and a preference for consistent and parsimonious information (Johnson-Laird 2010).
@@
Feasibility is one of the most agreed-upon metrics when discussing counterfactual explanations, although discussed under different names: Controllability (Byrne 2019), Actionability (Rasouli and Chieh Yu 2024) and split into Actionability and Mutability (Karimi et al. 2022). While actionability has been employed as a quantitative measure (Guidotti 2022), feasibility refers to whether the proposed changes are perceived as achievable and realistic. Explanations failing this criterion are rated poorly (Butz et al. 2024).
@@
Completeness previously discussed as Incompleteness (Zemla et al. 2017) or Informativeness, the latter of which also includes the notion of extraneous information (Stepin et al. 2022), is tied to understanding causal relations and partially depends on domain knowledge (Keil 2006). Evaluating completeness is challenging since people often fill logical gaps in explanations (Strickland and Keil 2011).
Finally, the dimension of Fairness in counterfactual explanations has been highlighted in recent work (Wang et al. 2024b). Concerns about models unintentionally encoding or amplifying biases in training data (Corbett-Davies et al. 2023) make it crucial to address potential unfairness and discrimination. Fairness is mostly been viewed as a quantitative metric (Ge et al. 2022), with limited understanding of its influence on perceived explanation quality.
@@
Relying on previous work on human preferences and explanatory virtues, we selected 8 different criteria capturing a range of relevant dimensions (see previous section for an overview) to guide the creation of diverse counterfactual scenarios. The scenarios in our study are grounded in actual outputs from counterfactual algorithms applied to commonly used datasets for counterfactual explanation evaluation, such as the Adult dataset and the Pima Indians Diabetes dataset. In some cases, we modified the algorithmic outputs to ensure a broader representation across explanatory quality metrics. Additionally, we tailored certain features to enhance clarity for human evaluators based on feedback from a pilot study. All counterfactual scenarios were designed from the perspective of improving the factual situation, as directionality has been shown to influence how explanations are perceived (Kuhl, Artelt, and Hammer 2023).
@@
The final version containing 30 counterfactual scenarios was shared on the Prolific platform and evaluated by 206 respondents. On average, completing the questionnaire took 42 minutes. All metrics were rated on scales detailed in Table 1, with a 6-point ordinal scale from 1 (lowest) to 6 (highest) except for Complexity, rated on a 5-point scale from -2 (too simple) to 2 (too complex), where 0 corresponded to desired complexity. Counterfactual scenarios were presented to the participants in randomised order, while the evaluation metrics remained in the same order. All respondents had to be at least 18 years of age and fluent in English to participate. The evaluated quationnaire is available on HuggingFace [^1] https://huggingface.co/datasets/anitera/CounterEval, and an example question is shown in Appendix A, Table A.1.
@@
After gathering and filtering questionnaire responses, further data processing was needed. The preprocessing scripts and model weights are publicly available in the GitHub repository [^2] https://github.com/anitera/CounterEval. For each question-metric pair, the average response from 196 participants was used as the final value. Complexity, originally rated on a -2 to 2 scale, was linearly scaled to align with the 1 to 6 scale used for other metrics. To minimize scale effects and enhance generalizability, we consolidated all metric values into three distinct categories: ”low,” ”medium,” and ”high.”. Data analysis suggested that the differences between scores of 1 and 2, 3 and 4, and 5 and 6 could be effectively compressed into these categories. This classification ensured a balanced distribution across the dataset. With 30 questions and 8 metrics per question, this resulted in 240 instances of metric evaluation in total.
@@
The optimal fine-tuning hyperparameters for every model were discerned through extensive testing (see Appendix C, Table C.1). All models were fine-tuned using a completion-only data collator from Huggingface’s trl library (Werra et al. 2020) to improve the predictive performance of the models. With a typical language modelling data collator, the model would have learned to predict the question text as well, but this was unnecessary for the task at hand. Due to its proprietary nature, GPT-4 was not used for fine-tuning.

### v3

@@ Abstract
As machine learning models evolve, maintaining transparency demands more human-centric explainable AI techniques. Counterfactual explanations, with roots in human reasoning, identify the minimal input changes needed to obtain a given output, and hence, are crucial for supporting decision-making. Despite their importance, the evaluation of these explanations often lacks grounding in user studies and remains fragmented, with existing metrics not fully capturing human perspectives. To address this challenge, we developed a diverse set of 30 counterfactual scenarios and collected ratings across 8 evaluation metrics from 206 respondents. Subsequently, we fine-tuned different Large Language Models (LLMs) to predict average or individual human judgment across these metrics. Our methodology allowed LLMs to achieve an accuracy of up to 63% in zero-shot evaluations and 85% (over a 3-classes prediction) with fine-tuning across all metrics. The fine-tuned models predicting human ratings offer better comparability and scalability in evaluating different counterfactual explanation frameworks.
@@ Introduction
The rapid adoption of AI in various domains has significantly increased the urgency for explainable AI models. Counterfactual explanations, which address the question ”How should the input be different in order to change the model’s decision outcome?” (Wachter, Mittelstadt, and Russell 2017), not only clarify the machine’s reasoning but also suggest potential changes that users might implement to achieve different results. These explanations enhance user trust and understanding by providing a richer mental representation compared to causal explanations (Warren, Byrne, and Keane 2023). Additionally, counterfactual explanations align closely with human cognitive processes (Miller 2019), as they provide alternative hypothetical realities that are pervasive in our natural reasoning (Byrne 2002).
@@
In light of these considerations, this paper addresses the following question: Can the evaluation process of counterfactual explanations be effectively automated using LLMs? To answer this question, we created a diverse set of 30 counterfactual scenarios that were designed to vary across multiple dimensions of explanatory qualities. The scenarios were evaluated by 206 human respondents in overall satisfaction, feasibility, consistency, completeness, trust, fairness, complexity, and understandability. Next, we divided data for fine-tuning several LLM models to assess every metric score and compared the results to human data on a reserved test set. The pipeline can be seen in Figure 1.
@@
  • First, we present a diverse dataset of human-evaluated counterfactual explanations CounterEval, encompassing a variety of metrics and scenarios, which could serve both for benchmarking and for training better causal representations of data, as demonstrated in (Chen et al. 2023).
@@
In recent years, some user studies have been conducted with tabular counterfactual data. For instance, (Warren, Byrne, and Keane 2023) conducted a study with 127 participants to compare the effects of counterfactual and causal explanations on objective prediction accuracy and subjective judgments of satisfaction and trust. (Bove et al. 2023) explored the impact of plural counterfactual examples on objective understanding and a modified Explanation Satisfaction Scale (Hoffman et al. 2018) in a lab study with 112 participants. (Förster et al. 2021) conducted a study with 46 participants assessing explanation realism and typicality. Two user studies have benchmarked counterfactual methods for perceived practicality of users in a study with 135 participants (Ghazimatin et al. 2020), and an online study with 500 responders (Spreitzer, Haned, and van der Linden 2022). Additionally, (Akula et al. 2022) tested their approach on image data, evaluating justified trust as a quantitative metric and explanation satisfaction as qualitative metric.
@@
Predicting human evaluation with machine learning has gained widespread acceptance in various domains, such as human-computer interaction (Kiseleva et al. 2016), recommendation systems (Siro, Aliannejadi, and De Rijke 2023), speech quality assessment (Reddy, Gopal, and Cutler 2022), etc. The advancement of LLMs’ causal reasoning abilities (Bhattacharjee et al. 2024) supports their use in explainability, as their natural language explanations exhibit qualities similar to human output (Castelnovo et al. 2024) and the explanatory process can be further enhanced through a post-output chat pipeline (Slack et al. 2023). LLMs have also been used to evaluate and model user satisfaction, providing insight into choices and preferences (Kim et al. 2024), and as artificial user-model tuning pairs (Gao et al. 2024).
@@
To select the dimensions for our study, we reviewed the literature on qualitative metrics that influence human judgments. Among the most frequently cited explanatory virtues are coherence and simplicity (Mackonis 2013), aligning with the understanding of human mental models and a preference for consistent and parsimonious information (Johnson-Laird 2010).
@@
Feasibility is one of the most agreed-upon metrics when discussing counterfactual explanations, although discussed under different names: Controllability (Byrne 2019), Actionability (Rasouli and Chieh Yu 2024) and split into Actionability and Mutability (Karimi et al. 2022). While actionability has been employed as a quantitative measure (Guidotti 2022), feasibility refers to whether the proposed changes are perceived as achievable and realistic. Explanations that fail this criterion are rated poorly (Butz et al. 2024).
@@
Completeness previously discussed as Incompleteness (Zemla et al. 2017) or Informativeness, the latter of which also includes the notion of extraneous information (Stepin et al. 2022), is tied to understanding causal relations and partially depends on domain knowledge (Keil 2006). Evaluating completeness is challenging, since people often fill logical gaps in explanations (Strickland and Keil 2011).
Finally, the dimension of Fairness in counterfactual explanations has been highlighted in recent work (Wang et al. 2024b). Concerns about models that unintentionally encode or amplify biases in training data (Corbett-Davies et al. 2023) make it crucial to address potential unfairness and discrimination. Fairness has been viewed mainly as a quantitative metric (Ge et al. 2022), with limited understanding of its influence on the perceived quality of the explanation.
@@
Relying on previous work on human preferences and explanatory virtues, we selected 8 different criteria capturing a range of relevant dimensions (see the previous section for an overview) to guide the creation of diverse counterfactual scenarios. The scenarios in our study are grounded in actual outputs from counterfactual algorithms applied to commonly used datasets for counterfactual explanation evaluation, such as the Adult dataset and the Pima Indians Diabetes dataset. In some cases, we modified the algorithmic outputs to ensure a broader representation across explanatory quality metrics. Additionally, we tailored certain features to enhance clarity for human evaluators based on feedback from a pilot study. All counterfactual scenarios were designed from the perspective of improving the factual situation, as directionality has been shown to influence the way explanations are perceived (Kuhl, Artelt, and Hammer 2023).
@@
The final version containing 30 counterfactual scenarios was shared on the Prolific platform and evaluated by 206 respondents. On average, completing the questionnaire took 42 minutes. All metrics were rated on scales detailed in Table 1, with a 6-point ordinal scale from 1 (lowest) to 6 (highest) except for Complexity, rated on a 5-point scale from -2 (too simple) to 2 (too complex), where 0 corresponded to desired complexity. Counterfactual scenarios were presented to participants in random order, while the evaluation metrics remained in the same order. All respondents had to be at least 18 years of age and fluent in English to participate. CounterEval dataset with annotated human responses is available on HuggingFace [^1] https://huggingface.co/datasets/anitera/CounterEval, and an example question is shown in Appendix A, Table A.1 (refer to the extended version of this paper for appendix (domnich2024unifyingevaluationcounterfactualexplanations)).
@@
After gathering and filtering questionnaire responses, further data processing was needed. The preprocessing scripts and model weights are publicly available in the GitHub repository [^2] https://github.com/anitera/CounterEval. For each question-metric pair, the average response from 196 participants was used as the final value. Complexity, originally rated on a -2 to 2 scale, was linearly scaled to align with the 1 to 6 scale used for other metrics. To minimize scale effects and enhance generalizability, we consolidated all metric values into three distinct categories: ”low,” ”medium,” and ”high”. Data analysis suggested that the differences between scores of 1 and 2, 3 and 4, and 5 and 6 could be effectively compressed into these categories. This classification ensured a balanced distribution across the dataset. With 30 questions and 8 metrics per question, this resulted in 240 instances of metric evaluation in total.
@@
The optimal fine-tuning hyperparameters for every model were discerned through extensive testing (see Appendix C, Table C.1). All models were fine-tuned using a completion-only data collator from Huggingface’s trl library (Werra et al. 2020) to improve the predictive performance of the models. With a typical language modeling data collator, the model would have learned to predict the question text as well, but this was unnecessary for the task at hand. Due to its proprietary nature, GPT-4 was not used for fine-tuning.